OpenStudy (anonymous):

solve for X 3x+32*14=454.9

OpenStudy (anonymous):

x would equal .1633 (repeating)

OpenStudy (anonymous):

That's not what I got

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

\[3x + 32 * 14 = 454.9\]

OpenStudy (anonymous):

The first thing we can do is simplify the multiplication

OpenStudy (anonymous):

so 32 * 14?

OpenStudy (anonymous):

448

OpenStudy (anonymous):

So now we have \[3x + 448 = 454.9\]

OpenStudy (anonymous):

What might be the next step if we wish to get x alone on one side

OpenStudy (anonymous):

3+448

OpenStudy (anonymous):

Well we can subtract 448 from both sides of the equation. What would this give us?

OpenStudy (anonymous):

\[3x+448 - 448 = 454.9 - 448\]

OpenStudy (anonymous):

6.9

OpenStudy (anonymous):

\[3x = 6.9\] What should be the last step? :)

OpenStudy (anonymous):

6.9 / 3

OpenStudy (anonymous):

so 2.3 :)

OpenStudy (anonymous):

thanks so much :)